The consumption of natural gas by a company satisfies
the empirical equation V = 1.50t + 0.008 00t ^2 where V
is the volume in millions of cubic feet and t the time in
months. Express this equation in units of cubic feet and
seconds. Put the proper units on the coefficients. Assume
a month is 30.0 days.

Grade:11

1 Answers

Ramesh V
14 years ago

V = at + bt2

V in million cubic feet

t in months

so after changing units to cubic feet and seconds

in a month, no. of seconds = 30*24*60*60 = 2.592 *106

so on substuting gives

V x 106 = 1.5*2.592 *106 t + 0.008*2.592 *106*2.592 *106*t2

V = 3.888 t + 53747.712 t2

with V in cubic feet and t in seconds

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch